OpenStudy (anonymous):

Jordan is a manager of a car dealership. He has two professional car washers, Matthew and Arianna, to clean the entire lot of cars. Matthew can wash all the cars in 14 hours. Arianna can wash all the cars in 11 hours. Jordan wants to know how long it will take them to wash all the cars in the lot if they work together. Write an equation and solve for the time it will take Matthew and Arianna to wash all the cars together. Explain each step.

OpenStudy (anonymous):

@surryyy @mathmath333 @help_people

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@acxbox22

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

HElp

OpenStudy (anonymous):

don't even know where to start!!

OpenStudy (kropot72):

In 1 hour Matthew, working alone, will wash 1/14 of the cars. In 1 hour Arianna, working alone, will wash 1/11 of the cars. Let the time taken to wash all of the cars working together be t hours. Then we can write: \[\large \frac{1}{14}+\frac{1}{11}=\frac{1}{t}\] @lizzieeej Can you follow this reasoning?

OpenStudy (anonymous):

Yes, but where from here? @kropot72

OpenStudy (kropot72):

Well, we have the required equation. Now we are asked to solve it to find the value of t.

OpenStudy (anonymous):

so add 1/14 and 1/11 ? @kropot72

OpenStudy (kropot72):

Yes, that is the first step.

OpenStudy (anonymous):

you get 25/154

OpenStudy (anonymous):

you can make it into decimal form if needed

OpenStudy (anonymous):

so 6.16

OpenStudy (kropot72):

Correct. So you now have \[\large \frac{25}{154}=\frac{1}{t}\]

OpenStudy (anonymous):

i made 25/154 equal 6.16 and then simplified to t=6.16

OpenStudy (kropot72):

Yes, the time taken is 6.16 hours.

OpenStudy (anonymous):

thanks so much!!!

OpenStudy (kropot72):

You're welcome :)